OpenStudy (anonymous):

Factor completely 64x2– 49

Parth (parthkohli):

Difference of squares: \( \color{Black}{\Rightarrow (8x)^2 - 7^2}\) a^2 - b^2 = (a + b)(a - b) can you do it now?

OpenStudy (unklerhaukus):

Use\[a^2b^2=(ab)^2\]

OpenStudy (anonymous):

Do you mean \[64x ^{2}-49\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[64=8^{2}\] Replace 64 with that, and you get: \[8^{2}x^{2}-7^{2}\] \[8^{2}x^{2}\] is the same as 8*8*x*x, or 8*x*8*x. That can also be written as (8*x)(8*x). That can be written as \[(8*x)^{2}\] From there follow what @ParthKohli posted.

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

(7x – 8)(7x + 8) (7x – 8)(7x – 8) (8x – 7)(8x – 7) (8x – 7)(8x + 7) These are the choices i have

OpenStudy (anonymous):

\[(8x)^{2}−7^{2}\] \[a^{2}−b^{2}\] so: \[(8x)^{2}=a^{2}\] \[7^{2}=b^{2}\] Known theorem: \[a^{2}−b^{2}=(a+b)(a−b)\]